• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name vmgray

    Log entry time 13:46:47 on December 3,2014

    Entry number 404378

    keyword=Target Alarm Handler Alarm Setting Changed

    After talking to JP about to the cryo temperature issue (see HALOG entries on target from last night and today), and that the cryo people will be working on the cryo system sometime today I have changed the limits on the alrams as such:

    minor high 16.1 K
    major high 16.5 K

    minor high 16.5 K
    major high 17.0 K

    For this one minor high was change eariler today (see HALOG entry 3310100)

    minor high 15.00 K
    major high 15.50 K

    minor high 15.50 K
    major high 16.00 K

    Both of these changes can be seen in the attachments as well.

    FIGURE 1

    FIGURE 2